### v1

@@ IX Conclusion and Limitations
We investigate the incorporation of uncertainty into speaker recognition systems and propose two uncertainty supervision strategies, together with a redesigned uncertainty estimation module. By integrating these components into a unified framework, termed $\mathcal{U}^{3}$-xi, we observe consistent and substantial performance improvements across multiple experimental settings. These results suggest that explicitly modeling uncertainty is a promising direction for advancing speaker recognition performance. Nevertheless, several limitations remain. In particular, uncertainty-aware scoring does not consistently improve minDCF, especially on cross-domain evaluation sets. This indicates that the uncertainty estimation for cross-domain data may not accurate enough. In future work, we plan to investigate more robust uncertainty estimation methods and examine the scalability of the proposed framework to large-scale training datasets.

### v2

@@ IX Conclusion and Limitations
We study uncertainty modeling in speaker recognition and propose two uncertainty supervision strategies together with a redesigned estimation module. These components are unified in the proposed $\mathcal{U}^{3}$-xi framework, which yields consistent and substantial performance gains across diverse experimental settings. The results demonstrate that explicitly modeling uncertainty is an effective and promising direction for improving speaker recognition.
Despite these gains, uncertainty-aware scoring does not consistently improve minDCF, particularly under cross-domain evaluation. This suggests that uncertainty estimation remains sensitive to domain mismatch and may not yet be sufficiently calibrated in such conditions. Future work will focus on more robust uncertainty modeling and on scaling the framework to larger and more diverse training corpora.